Home‎ > ‎Crew 162‎ > ‎

0130-GreenHabReport

GreenHab Report
01/30/2016
Renee Garifi

Current Temperature:
70F/21C


Crew 161 Plant Report:
The radish and bean sprouts that were transferred and repotted are doing
alright. The best growing plants are the ones that were originally
sprouted in 100% Earth soil with fertilizer. Sprouts that died before
they could be transferred included the seedlings grown in 100% Mars soil
without fertilizer. Photos to follow of the transferred sprouts. Added
60mL of water to each set of transferred sprouts.

OASYS Passive Watering System:
Seed wicks remain moist. No cabbage sprouts yet. Added 1L of water to
reservoir on 1/29 and 0.5L on 1/30.

Active Watering System:
Added 1L of water to reservoir on 1/29 and 1L on 1/30. No cabbage
sprouts yet but wicks remain moist.

Confirmed power strip was on and functional. Will ask PI about whether
lights should be cycled and how to get more light to the Active Watering
System on the lower shelf of the temporary grow tent.